OpenStudy (anonymous):

What mass of Copper (II) nitrate would be produced from the complete reaction of 45.6 grams of copper according to the chemical reaction shown below

OpenStudy (aaronq):

for problems involving stoichiometry, you have to convert whatever information they give you to moles (signified by the symbol "n") Use the stoichiometric coefficients to build a ratio, like this: e.g. for a general reaction: \(\color{red}{a}A + \color{blue}{b}B\) \(\rightleftharpoons\) \( \color{green}{c}C\) where upper case are the species (A,B,C), and lower case (a,b,c) are the coefficients (in red), \(\dfrac{n_A}{\color{red}{a}}=\dfrac{n_B}{\color{blue}{b}}=\dfrac{n_C}{\color{green}{c}}\) from here, you can solve for moles then convert to grams. To convert mass to moles, use the relationship: \(n=\dfrac{m}{M}\) where, M=molar mass, m=mass, and n= moles.

OpenStudy (anonymous):

\[Cu+2AgNO_3\rightarrow Cu(NO_3)_2 +2Ag\]